OpenStudy (anonymous):

solve the system of equations below. 2x + 2y = 12 3x + 2y = 17

OpenStudy (anonymous):

do the same thing you just did with pooja but subtract 2-3=? 2-2=? 12-17=?

OpenStudy (anonymous):

i got 1x + 0y = 5 but what do i do next?

OpenStudy (anonymous):

it actually -x=-5 but that equals x=5... anyways. now you substitute x in one of the equations (lets use the first one) so 2(5)+2y=12

OpenStudy (anonymous):

10+2y=12 -10 -10 2y=2 y=1 and x=5